OpenStudy (anonymous):

The length of a rectangle is twice the width. The area is 50yd^2.What is the length and width? The length is __?__yd The width is ___?__yd

OpenStudy (saifoo.khan):

Area = l * b Area = 2w(w) 50 = 2w^2 w^2 = 50. w = 5. Therefore Length = 10 Width = 5

OpenStudy (anonymous):

thanks :)

OpenStudy (saifoo.khan):

No Problem.